# Machine-directed gravitational-wave counterpart discovery

### v2

@@ 4.1 Human benchmark
We solicited the aid of six volunteer astronomers to solve our KN follow-up problem. Similar to Pythia they completed a training and testing phase, the former with as many episodes are they chose and the latter with 100 episodes. During both phases, we recorded their choices and decision times. Finally, we asked them to self-identify as experts and non-experts in KN searches and provide a short description of their strategies (summarized in Appendix A).
@@ 4.2 Agent preference/policy interpretation
The first point of difference is the agent’s behavior when they perform poorly. In such cases, Pythia allocates the bulk of the budget to a single event (see episode P3 and P4 in Figure 2). On the other hand, humans spread their budget (see episode H3 in Figure 3), sampling several events to try to rule them out (the intent was communicated to us in discussions after the test). The AI does not explore in failure cases simply because it is not rewarded for doing so. During training, when it makes the wrong choice, the weights do not get updated by much, so it is biased by the weight initialization/experience thus far. In these cases, not updating the target event might be an effective strategy as in high reward/success cases the agent needs to stick with the same event. This aspect underscores the importance of carefully formulating the reward in design of artificial agents.
@@ 5 Conclusions
In this paper we develop the first AI agent capable of strategizing a sequential transient follow-up program. Given several events with evolving light curves, one of which is a KN and the rest are contaminants, the agent must maximize follow-up for true candidate, using information learned from actions taken and routine survey observations to improve its strategy in the next round. Using a simple behavior policy (linear VFA) our agent demonstrates that the problem is learnable by machines and comes close to human performance. However, the agent is quite greedy in its decisions (the score does not improve much with time) suggesting that the agent’s hypothesis function is an insufficient representation of the optimal behavior policy. More complex agents (e.g. using deep $Q$ networks or policy gradient algorithms) could help bridge the gap with human experts. This could really help streamline KN search efforts. In the development of such agents it would also be useful to employ graph neural networks to represent the light curves as this is a more natural choice to ensure learning order invariant policies than the convolutional neural networks used here.
@@ Appendix A Strategies of human agents
The agent’s initial choice was random between any red transient. They later focussed on events with fade rates that were neither too fast (indicative of GRB afterglows) nor too slow (indicative of normal SNe). They were interested in events with an unusual color evolution. They also preferred late observations of single detection events to rule them out.

### v3

@@ 4.1 Human benchmark
We solicited the aid of six volunteer astronomers to solve our KN follow-up problem. Similar to Pythia they completed a training and testing phase, the former with as many episodes as they chose and the latter with 100 episodes. During both phases, we recorded their choices and decision times. Finally, we asked them to self-identify as experts and non-experts in KN searches and provide a short description of their strategies (summarized in Appendix A).
@@ 4.2 Agent preference/policy interpretation
The first point of difference is the agent’s behavior when they perform poorly. In such cases, Pythia allocates the bulk of the budget to a single event (see episode P3 and P4 in Figure 2). On the other hand, humans spread their budget (see episode H3 in Figure 3), sampling several events to try to rule them out (the intent was communicated to us in discussions after the test). The AI does not explore in failure cases simply because it is not rewarded for doing so. During training, when it makes the wrong choice, the weights do not get updated by much, so it is biased by the weight initialization/experience thus far. In these cases, not updating the target event might be an effective strategy as in high reward/success cases the agent needs to stick with the same event. This aspect underscores the importance of carefully formulating the reward in the design of artificial agents.
@@ 5 Conclusions
In this paper we develop the first AI agent capable of strategizing a sequential transient follow-up program. Given several events with evolving light curves, one of which is a KN and the rest are contaminants, the agent must maximize follow-up for the true candidate, using information learned from actions taken and routine survey observations to improve its strategy in the next round. Using a simple behavior policy (linear VFA) our agent demonstrates that the problem is learnable by machines and comes close to human performance. However, the agent is quite greedy in its decisions (the score does not improve much with time) suggesting that the agent’s hypothesis function is an insufficient representation of the optimal behavior policy. More complex agents (e.g. using deep $Q$ networks or policy gradient algorithms) could help bridge the gap with human experts. This could really help streamline KN search efforts. In the development of such agents it would also be useful to employ graph neural networks to represent the light curves as this is a more natural choice to ensure learning order invariant policies than the convolutional neural networks used here.
@@ Appendix A Strategies of human agents
The agent’s initial choice was random between any red transient. They later focused on events with fade rates that were neither too fast (indicative of GRB afterglows) nor too slow (indicative of normal SNe). They were interested in events with an unusual color evolution. They also preferred late observations of single detection events to rule them out.